OpenStudy (anonymous):

The base of a rectangular solid is a square with a side of the length a cm . The height of the rectangular solid is b cm and its volume is equal to V cm^3. Express the variable a in terms of b and V.

jhonyy9 (jhonyy9):

V=a*a*b --- a^2 = V/b --- a=sqrt(V/b)

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

ok